# Designing Human-AI Systems: Anthropomorphism and Framing Bias on Human-AI Collaboration

### v1

@@ 3.4 Experimental design and hypotheses
The experiment followed a factorial design of two variables and three treatment levels \@BBOPcitep\@BAP\@BBN(Kirk,, 2009)\@BBCP. Once a subject started the survey, a specific number from one to nine was randomly assigned to them. Each number had a corresponding treatment combination assigned and remained the same for participants during the whole survey flow. The dependent variable for the two hypotheses was the correlation between the AI recommendation and the subject’s decision, indicating the frequency of agreement or disagreement between the two.
@@ 3.4.1 Hypothesis 1 (H1)
Note that equivalent frames could have been achieved with different descriptors: I could have labeled the positive outcome with a negative descriptor (“80% probability of not being a failure”), and the negative outcome with a positive descriptor (“20% probability of not being a success”). However, unlike outcome-framing manipulation, descriptors do not seem to have a significant impact on attribute-framing bias. Therefore, I opted for a balanced presentation of both positive and negative framings in the neutral treatment level to minimize the framing effect, as prior work on the subject suggests \@BBOPcitep\@BAP\@BBN(Kreiner and Gamliel,, 2019)\@BBCP. To further ensure neutrality in the neutral treatment level, descriptors, and the estimated probabilities were displayed in a table. Moreover, I designed two variants of the table with distinct frame order: one table had the positively framed information in the first column (Figure 4), and the other table placed the same information in the second column (Figure 5). Subjects who were assigned to a treatment combination with neutral framing were randomly assigned to one of these variations.
@@
To verify frames’ valence, I corroborated the polarity of the words “success” and “failure” with the Subjectivity Lexicon presented in the work “Recognizing Contextual Polarity in Phrase-Level Sentiment Analysis” \@BBOPcitep\@BAP\@BBN(Wilson et al.,, 2005)\@BBCP. According to the Lexicon, these words display a positive and negative polarity out of context, and these polarities are considered weakly subjective.
@@ 3.4.2 Hypothesis 2 (H2)
In the second hypothesis (H2) I wanted to measure the impact of AI’s anthropomorphic design on human trust. To test the hypothesis, I designed three different levels of anthropomorphism, or AI identities, which differed in the nature of the features attributed to the AI (human-like, robot-like, and generic). In the human-like condition, the AI was given humanlike features: it was referred to by a human name (Maya), was given a gender (female), and was given a human appearance (using an image of a real human). In the robot-like condition, the AI had a robotic name (ZAX-D2), an impersonal pronoun (it), and was represented with a humanoid robot image. For the generic condition (the control group), I omitted any reference to its identity by replacing the names given in the robot-like and human-like conditions with the word “AI”. In addition, no picture was provided.
@@
It should be mentioned that there are many ways in which human-like and robot-like features can be attributed to AI. I am aware of the fact that more complex cues could be introduced to enhance the anthropomorphic effect. For instance, \@BBOPcitet\@BAP\@BBNSchanke et al., (2021)\@BBCP used humor, social presence, and communication delays to induce anthropomorphism in the interaction with chatbots. However, these cues might require the implementation of a more sophisticated interface.
It is still unclear the number and type of human-like characteristics needed to cause anthropomorphism \@BBOPcitep\@BAP\@BBN(De Visser et al.,, 2016)\@BBCP. I followed a similar anthropomorphic design to the one proposed by \@BBOPcitet\@BAP\@BBNPak et al., (2012)\@BBCP, which reported the influence of human-like clues in subjects’ trust through the use of images. Additionally, I integrated subtle linguistic cues, such as the use of pronouns (e.g., “she” vs “it”).
@@ 3.5 Results
To investigate the impact of anthropomorphism and framing on subjects’ agreement with AI recommendations, four t-tests were designed, each comparing a treatment group with its respective control group (positive-neutral framing, negative-neutral framing, human-generic identity, and robot-generic identity). These t-tests were run three times, using different sample sizes based on attention check performance. The first iteration included all subjects exposed to a specific treatment condition, regardless of the number of attention checks passed ($a\geq 0$). The second iteration considered subjects that passed at least one attention check ($a\geq 1$). Finally, the third iteration included only subjects that passed the two attention checks ($a=2$). As shown in Table 1, a considerable percentage of subjects did not pass at least one attention check.
Identity Framing
@@
% of subjects failing both 10% 8% 12% 15% 7% 7%
% of subjects failing at least 1 49% 51% 55% 58% 45% 54%
Table 1: Subjects’ attention check performance by treatment group. About half of the 749 subjects that finished the survey did not pass both attention checks.
I assumed that subjects agreed with the AI recommendation if they decided to invite a candidate that had a probability of success greater than or equal to 50%, according to the AI model; or if they did not invite a candidate that had a probability of success lower than 50%. Therefore, to calculate agreement rates, I counted the number of times a subject agreed with the AI recommendation and divided it by the total number of questions.
Table 2 shows the results of the t-test analysis. It displays the difference in agreement rates’ means, the t statistic, the p-values, and their corresponding significance. If the treated group has a greater rate of agreement with the AI than the control group, the difference is positive; whereas if the treated group has a lower rate of agreement with the AI than the control group, the difference has a negative sign.
Framing Anthropomorphism
Positive Negative Human Robot
$a\geq 0$ $a\geq 1$ $a=2$ $a\geq 0$ $a\geq 1$ $a=2$ $a\geq 0$ $a\geq 1$ $a=2$ $a\geq 0$ $a\geq 1$ $a=2$
mean diff 0.012 0 -0.009 -0.01 -0.012 -0.019 -0.019 -0.024 -0.04 -0.023 -0.017 -0.036
t-test -0.852 -0.016 0.452 0.691 0.866 1.022 1.343 1.677 2.077 1.601 1.145 1.81
df 489 433 211 499 465 239 498 443 229 499 450 231
p-value 0.395 0.987 0.652 0.49 0.387 0.308 0.18 0.094 0.039 0.11 0.253 0.072
significance - - - - - - - + * - - +
Table 2: Summary of the results from the t-tests run between the treatment groups and their corresponding control group. Each of the twelve columns represents a distinct t-test. These columns display the observed difference in the rates of agreement between groups, along with their statistical significance. For each treatment condition, three different t-tests were conducted, differing on the number of attention checks passed by subjects.
Regarding the framing bias effect, the results of the study did not support the hypothesized relationship between framing and agreement rates. No statistically significant differences in agreement rates were found between the positive and negative framing conditions, regardless of attention check performance. This held true across all three repetitions of the t-tests. The p-values associated with these findings were consistently above 0.1, indicating a lack of significant difference between the treatment and control groups.
In relation to anthropomorphism and the effects of AI identities, no significant difference was found by examining the data from all available subjects. Nevertheless, if we only analyze subjects that passed at least one attention check, we discover that subjects exposed to the human condition were on average 2.4% less likely to agree with the AI in any given question compared to the control group, with a significance level of $p<0.1$, and $t(443)=1.677$. The difference in agreement rates between the human condition and the control group becomes even more pronounced (4%), and significant $(p<0.05,t(229)=2.077)$, when considering subjects who completed both attention checks successfully. Likewise, for this subset, subjects exposed to the robot condition were 3.6% less likely to agree with the AI, with a significance level of $p<0.1$ and $t(231)=1.81$.
The study’s findings indicate that anthropomorphism has a significant impact on agreement rates. However, contrary to expectations, the results show that subjects are more likely to agree with the AI’s recommendations when it does not have a human or robot identity (as shown in Figure 8).
Figure 8: Confidence intervals for rates of agreement by identity treatment, with a 90% confidence level. Only subjects who successfully passed the two attention checks were included. There is a 90% chance that the mean of the whole population falls somewhere between 0.695 and 0.719 for the generic condition; 0.651 and 0.674 for the robot condition; and 0.649 and 0.67 for the human condition.
@@
In sum, none of the stated hypotheses has been validated by the experiment. The framing of the AI recommendations did not significantly impact the agreement rate of subjects. On the other hand, there was a statistically significant difference in agreement rates caused by the identity of the AI, with the human-like and robot-like AI identities leading to lower agreement rates compared to the control group. However, this finding contradicts the initial expectation that a human-like AI would result in higher agreement.
@@ 4 Discussion
In this study, I analyzed the impact of cognitive biases on human-AI collaboration by conducting an experiment that replicated the recruitment process of a company. The experiment aimed to investigate the effect of framing bias and anthropomorphism on hiring managers’ decisions when they are assisted by AI recommendations. By gaining a deeper understanding of these biases, and their impact on human-AI collaboration, my goal was to provide valuable insights that can inform the development of AI-based tools, and ultimately enhance the synergy between humans and AI in decision-making processes.
@@ 4.1 Framing effect
Although attribute framing has been extensively demonstrated in the literature \@BBOPcitep\@BAP\@BBN(Freling et al.,, 2014)\@BBCP through numerous studies \@BBOPcitep\@BAP\@BBN(Davis and Bobko,, 1986; Duchon et al.,, 1989; Dunegan,, 1993; Leong et al.,, 2017; Levin, 1987a, )\@BBCP, the framing of AI recommendations did not significantly affect subjects’ agreement rate in this study. A possible explanation could be that the manipulation chosen for the experiment was insufficient to influence subjects in their decisions. While previous studies were able to induce subjects into framing bias through success and failure rates, the manipulation differed in certain aspects.
@@
For instance, participants in this study received additional information about candidates’ education and work experience alongside the AI recommendations. In contrast, other studies that have demonstrated framing effects provided minimal information to subjects. In Levin’s experiment \@BBOPciteyear\@BAP\@BBN\@BBCP, subjects evaluating basketball players only had access to the percentage of shots missed or passed. Likewise, the study of \@BBOPcitet\@BAP\@BBNDuchon et al., (1989)\@BBCP simulated an R&D financial allocation decision where the main information available to subjects was the rate of success or failure of previous projects. The discrepancy between the experimental design of these studies and my own suggests that providing additional information, such as the candidate’s education and work experience, may have a debiasing effect. This could potentially explain the absence of a significant framing effect in my study.
@@
The hypothetical debiasing effect of additional information in our experiment is consistent with Kahneman’s dual-system theory, specifically, with the “What You See Is All There Is” principle which rules System 1. This principle suggests that people tend to focus on the information that is presented to them. In the case of framing effects, they can be explained by the way that the negative and positive frames capture individuals’ attention, leading them to overlook the implicit counterparts \@BBOPcitep\@BAP\@BBN(Kahneman,, 2011)\@BBCP. Although a rate of success equal to 90% implies a rate of failure equal to 10% (and vice-versa), this equivalence is not immediately apparent to individuals. Building on this concept, \@BBOPcitet\@BAP\@BBNKreiner and Gamliel, (2018)\@BBCP successfully moderated framing bias by shifting subjects’ attention to the complementary, non-explicit frame. A similar phenomenon may take place in my experiment. Additional information may have redirected subjects’ attention away from the initially presented frame in the AI recommendation, thereby mitigating the framing effect and debiasing subjects’ judgments.
@@
Since currently humans rarely rely exclusively on AI decision-aid tools for decision-making, this finding might suggest that the framing effect is not a significant factor that interferes with human-AI collaboration for the moment, provided that the decision-makers have additional information on which to base their judgments. Indeed, some research has found no evidence of the framing effect impacting the reliance of humans on decision-aid systems. For instance, in Brown and Jones’ experiment \@BBOPciteyear\@BAP\@BBN1998\@BBCP, subjects had to decide which computer system to recommend to their clients based on the scores of some features. Additionally, they were assisted by a decision-aid algorithm, which computed those scores and recommended a specific computer system to subjects. Subjects were told that they could potentially gain or lose an important client. However, the frame of the scenario did not significantly affect the reliance of subjects on the decision-aid system. A different experiment conducted by \@BBOPcitep\@BAP\@BBN(Huerta et al.,, 2012)\@BBCP on assisted decision-making suggests that cultural factors may determine the effects of framing on decision-aid reliance. For the American sample, the framing of a report generated by an automated system did not influence the intention to investigate fraud, whereas it did influence the Mexican sample.
@@
The lack of studies that investigate framing effects in the context of AI decision-aid systems points to the need for a deeper understanding of the phenomenon. Although framing effect may not be a pressing concern currently given the limited scope of AI tools, the rapid advancements in AI technologies might lead to increased reliance on AI-driven decision-making in the future. In this experiment, the AI recommendation was one of the several factors presented to candidates. The additional information provided may have shifted subjects’ attention from the frame, mitigating bias. However, it is possible to imagine a future where AI systems are designed to omit external information, or to display solely AI-generated information, potentially introducing framing bias.
@@ 4.2 Anthropomorphism and AI identities
The findings of this study reveal a significant effect of AI identities on agreement among subjects who successfully completed the two attention checks of the survey. Nevertheless, the observed effect was not expected. AI identities were found to have a negative impact on agreement with AI recommendations, regardless of their level of anthropomorphism. Subjects exposed to both human-like and robot-like AI identities had a significantly lower agreement rate in comparison to the control group. Consequently, the results of the experiment deviate from the results of the reference paper, where a human image increased reliance on automated decision-aid systems \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP.
@@
One potential explanation for the divergence in results may lie in the differences between both manipulations. In contrast to the reference study, which used an image of a person dressed as a doctor, the generic human image used in my experiment did not convey a specific professional affiliation, such as someone working in a company’s HR department. Moreover, the structure of the recommendation also differed between studies. While my study’s recommendation was presented in terms of probabilities, the other study employed what could be considered a more human-like structure (e.g., “my advice is: cholesterol”). This potential explanation would imply that in order to increase agreement with an AI tool, the credibility of AI’s human identity is something that should be considered, especially regarding how it conveys information to the decision-maker.
@@
Another potential explanation for the observed effect of AI identities on subjects’ agreement with the recommendation, and its deviation from the reference paper, is the different nature of the interaction. In the case of a decision-aid system that assists patients in making diabetes management decisions, the system is replacing a human in an already existent and social interaction. It is common to go to the doctor to get assistance on how to better manage our health. On the contrary, in my study, the AI had a task-oriented interaction with subjects in a nonsocial setting. As \@BBOPcitet\@BAP\@BBNYang et al., (2022)\@BBCP argue, when the interaction occurs in nonsocial scenarios, AI products are treated as utilities, resulting in an indifference toward its degree of anthropomorphism. Consequently, contrary to the established belief that human-like AI identities facilitate human-AI interaction, in certain cases, designers of AI tools should not focus their efforts on anthropomorphism \@BBOPcitep\@BAP\@BBN(Yang et al.,, 2022)\@BBCP. The findings of my experiment align with this explanation, suggesting that a more neutral identity would be preferable in certain cases.

### v2

@@ 3.4 Experimental design and hypotheses
The experiment followed a factorial design of two variables and three treatment levels \@BBOPcitep\@BAP\@BBN(Kirk,, 2009)\@BBCP. It had a 3 x 3 between-subjects design, with three different AI identities (human-like, robot-like, generic) and three types of framing (positive, negative, and neutral). Once a subject started the survey, a specific number from one to nine was randomly assigned to them. Each number had a corresponding treatment combination assigned and remained the same for participants during the whole survey flow. The dependent variable for the two hypotheses was the correlation between the AI recommendation and the subject’s decision, indicating the frequency of agreement or disagreement between the two.
@@ 3.4.1 Hypothesis 1 (H1)
Note that equivalent frames could have been achieved with different descriptors: we could have labeled the positive outcome with a negative descriptor (“80% probability of not being a failure”), and the negative outcome with a positive descriptor (“20% probability of not being a success”). However, unlike outcome-framing manipulation, descriptors do not seem to have a significant impact on attribute-framing bias. Therefore, we opted for a balanced presentation of both positive and negative framings in the neutral treatment level to minimize the framing effect, as prior work on the subject suggests \@BBOPcitep\@BAP\@BBN(Kreiner and Gamliel,, 2019)\@BBCP. To further ensure neutrality in the neutral treatment level, descriptors, and the estimated probabilities were displayed in a table. Moreover, we designed two variants of the table with distinct frame order: one table had the positively framed information in the first column (Figure 4), and the other table placed the same information in the second column (Figure 5). To neutralize unintended effects due to the ordering of the columns, subjects exposed to the neutral framing were randomly assigned to one of these two variations.
@@
To verify frames’ valence, we corroborated the polarity of the words “success” and “failure” with the Subjectivity Lexicon presented in the work “Recognizing Contextual Polarity in Phrase-Level Sentiment Analysis” \@BBOPcitep\@BAP\@BBN(Wilson et al.,, 2005)\@BBCP. According to the Lexicon, these words display a positive and negative polarity out of context, and these polarities are considered weakly subjective.
@@ 3.4.2 Hypothesis 2 (H2)
In the second hypothesis (H2) we wanted to measure the impact of AI’s anthropomorphic design on human trust. To test the hypothesis, we designed three different AI identities, which differed in the nature of the features attributed to the AI (human-like, robot-like, and generic). In the human-like condition, the AI was given humanlike features: it was referred to by a human name (Maya), was given a gender (female), and was given a human appearance (using an image of a real human). In the robot-like condition, the AI had a robotic name (ZAX-D2), an impersonal pronoun (it), and was represented with a humanoid robot image. For the generic condition (the control group), we omitted any reference to its identity by replacing the names given in the robot-like and human-like conditions with the word “AI”. In addition, no picture was provided.
@@
It should be mentioned that there are many ways in which human-like and robot-like features can be attributed to AI. We are aware of the fact that more complex cues could be introduced to enhance the anthropomorphic effect. For instance, \@BBOPcitet\@BAP\@BBNSchanke et al., (2021)\@BBCP used humor, social presence, and communication delays to induce anthropomorphism in the interaction with chatbots. However, these cues might require the implementation of a more sophisticated interface.
The specific number and nature of human-like characteristics required to elicit perceptions of a system as humanlike remain unclear \@BBOPcitep\@BAP\@BBN(De Visser et al.,, 2016)\@BBCP. In response to this uncertainty, we adopted an anthropomorphic design similar to the one proposed by \@BBOPcitet\@BAP\@BBNPak et al., (2012)\@BBCP, which reported the influence of human-like clues in subjects’ trust through the use of images. Additionally, we integrated subtle linguistic cues, such as the use of pronouns (e.g., “she” vs “it”).
@@ 3.5 Results
To investigate differences in subjects’ agreement rates with AI recommendations as a function of framing and AI identity, a 3 (framing: positive, negative, neutral) X 3 (AI identity: human-like, robot-like, generic) ANOVA with unbalanced replications was run. Since a large interaction between the two factors was not expected, we used a Type II ANOVA \@BBOPcitep\@BAP\@BBN(Smith and Cribbie,, 2021)\@BBCP. We assumed subjects agreed with the AI if they invited candidates to which the AI assigned a probability of success greater than or equal to 50% or if they declined to invite a candidate to which the AI assigned a probability of success smaller than 50%. To calculate agreement rates, we counted the number of times a subject agreed with the AI and divided it by the total number of questions.
As shown in Table 1, a considerable percentage of subjects failed at least one attention check. Dropping subjects who failed these attention checks could potentially cause some unintended biases \@BBOPcitep\@BAP\@BBN(Aronow et al.,, 2019)\@BBCP. To mitigate this effect, we decided to run the ANOVA three times using different sample sizes based on attention check performance. The first ANOVA iteration included all subjects exposed to a specific treatment condition, regardless of the number of attention checks passed ($a\geq 0$). The second iteration considered subjects that passed at least one attention check ($a\geq 1$). Finally, the third iteration included only subjects who passed the two attention checks ($a=2$).
AI Identity Framing
@@
Table 1: Subjects’ attention check performance by treatment group.
[…]
Framing AI Identity Interaction
$a\geq 0$ $a\geq 1$ $a=2$ $a\geq 0$ $a\geq 1$ $a=2$ $a\geq 0$ $a\geq 1$ $a=2$
F-statistic 1.183 0.481 0.504 1.488 1.448 2.554 1.186 1.176 1.547
p-value 0.307 0.618 0.604 0.226 0.236 0.079 0.316 0.32 0.188
significance - - - - - + - - -
Table 2: Summary of the Type II ANOVA results for the two different factors (AI identity, Framing) and their interactions. The ANOVA was run three times with different sample sizes, categorized by attention check performance. The table includes the F-statistics, the p-values, and their corresponding significance levels. Treatments with a p-value greater than 0.1 are indicated by a hyphen (”-”), whereas treatments with a p-value less than 0.1 are indicated by a plus sign (”+”).
To further investigate the effect of AI identities on subjects who passed both attention checks, we conducted a post-hoc analysis using the Tukey-Kramer method to account for the unbalanced design of the experiment. As shown in Figure 8, there is a significant difference in agreement rates between the generic and human identity conditions. Subjects exposed to the human condition were on average 3.99% less likely to agree with the AI in any given question compared to the control group, with a significance level of $p=0.0987$. The study’s findings indicate that attributing a human-like identity to the AI significantly impacts agreement rates. However, contrary to our expectations, the results show that subjects are more likely to agree with the AI’s recommendations when it lacks a human identity.
0.080.08-0.08- 0.080.040.04-0.04- 0.04000.040.040.040.040.080.080.080.08Human - RobotGeneric - RobotGeneric - HumanMean Difference
Figure 8: Mean differences between AI identities with 90% confidence intervals based on Tukey HSD test. The dashed vertical line represents zero difference. Intervals that do not cross this line indicate significant differences ($p<0.10$).
@@
In sum, none of the stated hypotheses has been validated by the experiment. The framing of the AI recommendations did not significantly impact the agreement rate of subjects. On the other hand, there was a statistically significant difference in agreement rates caused by the identity of the AI, with the human-like AI identity leading to lower agreement rates compared to the control group. This finding contradicts the initial expectation that a human-like AI would result in higher agreement.
@@ 4 Discussion
In this study, we analyzed the impact of cognitive biases on human-AI collaboration by conducting an experiment that replicated the recruitment process of a company. The experiment aimed to investigate the effect of framing bias and anthropomorphism on hiring managers’ decisions when they are assisted by AI recommendations. By gaining a deeper understanding of these biases, and their impact on human-AI collaboration, our goal was to provide valuable insights that can inform the development of AI-based tools, and ultimately enhance the synergy between humans and AI systems in decision-making processes.
@@ 4.1 Framing effect
Although attribute framing has been extensively demonstrated in the literature \@BBOPcitep\@BAP\@BBN(Freling et al.,, 2014)\@BBCP through numerous studies \@BBOPcitep\@BAP\@BBN(Davis and Bobko,, 1986; Duchon et al.,, 1989; Dunegan,, 1993; Leong et al.,, 2017; Levin, 1987a, )\@BBCP, the framing of AI recommendations did not significantly affect subjects’ agreement rate in our experiment. A possible explanation could be that the manipulation chosen for the experiment was insufficient to influence subjects in their decisions. While previous studies were able to induce subjects into framing bias through success and failure rates, the manipulation differed in certain aspects.
@@
For instance, participants in this study received additional information about candidates’ education and work experience alongside the AI recommendations. In contrast, other studies that have demonstrated framing effects provided minimal information to subjects. In Levin’s experiment \@BBOPciteyear\@BAP\@BBN\@BBCP, subjects evaluating basketball players only had access to the percentage of shots missed or passed. Likewise, the study of \@BBOPcitet\@BAP\@BBNDuchon et al., (1989)\@BBCP simulated an R&D financial allocation decision where the main information available to subjects was the rate of success or failure of previous projects. The discrepancy between the experimental design of these studies and our own suggests that providing additional information, such as the candidate’s education and work experience, may have a debiasing effect. This could potentially explain the absence of a significant framing effect in our study.
@@
The hypothetical debiasing effect of additional information in our experiment is consistent with Kahneman’s dual-system theory, specifically, with the “What You See Is All There Is” principle which rules System 1. This principle suggests that people tend to focus on the information that is presented to them. In the case of framing effects, they can be explained by the way that the negative and positive frames capture individuals’ attention, leading them to overlook the implicit counterparts \@BBOPcitep\@BAP\@BBN(Kahneman,, 2011)\@BBCP. Although a rate of success equal to 90% implies a rate of failure equal to 10% (and vice-versa), this equivalence is not immediately apparent to individuals. Building on this concept, \@BBOPcitet\@BAP\@BBNKreiner and Gamliel, (2018)\@BBCP successfully moderated framing bias by shifting subjects’ attention to the complementary, non-explicit frame. A similar phenomenon may take place in our experiment. Additional information may have redirected subjects’ attention away from the initially presented frame in the AI recommendation, thereby mitigating the framing effect and debiasing subjects’ judgments.
@@
Since currently humans rarely rely exclusively on AI decision-aid tools for decision-making, this finding might suggest that the framing effect is not a significant factor that interferes with human-AI collaboration for the moment, provided that the decision-makers have additional information on which to base their judgments. Indeed, some research has found no evidence of the framing effect impacting the reliance of humans on decision-aid systems. For instance, in Brown and Jones’ (\@BBOPciteyear\@BAP\@BBN1998\@BBCP) experiment, subjects had to decide which computer system to recommend to their clients based on the scores of some features. Additionally, they were assisted by a decision-aid algorithm, which computed those scores and recommended a specific computer system to subjects. Subjects were told that they could potentially gain or lose an important client. However, the frame of the scenario did not significantly affect the reliance of subjects on the decision-aid system. A different experiment conducted by \@BBOPcitep\@BAP\@BBN(Huerta et al.,, 2012)\@BBCP on assisted decision-making suggests that cultural factors may determine the effects of framing on decision-aid reliance. For the American sample, the framing of a report generated by an automated system did not influence the intention to investigate fraud, whereas it did influence the Mexican sample.
@@
Although AI systems are increasingly present in our society, the framing effect has not yet emerged as a major concern. The lack of studies on this topic serves as proof. However, as AI applications continue to expand, the potential effects of framing bias in AI-assisted decision-making could become increasingly relevant. In this experiment, the AI recommendation was one of the several factors presented to candidates. The additional information provided may have shifted subjects’ attention from the frame, mitigating bias. Nevertheless, it is possible to imagine a future where AI systems are designed to omit external information, or to display solely AI-generated information, potentially introducing framing bias.
@@ 4.2 Anthropomorphism and AI identities
The findings of this study reveal a significant effect of the human AI identity on agreement rates among subjects who successfully completed the two attention checks of the survey. Nevertheless, the observed effect was not expected. The human identity was found to have a negative impact on agreement with AI recommendations. Subjects exposed to this identity had a significantly lower agreement rate in comparison to the control group. Consequently, the results of the experiment deviate from the results of the reference paper, where a human image increased reliance on automated decision-aid systems \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP.
@@
One potential explanation for the divergence in results may lie in the differences between both manipulations. In contrast to the reference study, which used an image of a person dressed as a doctor, the generic human image used in our experiment did not convey a specific professional affiliation, such as someone working in a company’s HR department. Moreover, the structure of the recommendation also differed between studies. While our study’s recommendation was presented in terms of probabilities, the other study employed what could be considered a more human-like structure (e.g., “my advice is: cholesterol”). This potential explanation would imply that in order to increase agreement with an AI tool, the credibility of AI’s human identity is something that should be considered, especially regarding how it conveys information to the decision-maker.
@@
Another potential explanation for the observed effect of AI identities on subjects’ agreement with the recommendation, and its deviation from the reference paper, is the different nature of the interaction. In the case of a decision-aid system that assists patients in making diabetes management decisions, the system is replacing a human in an already existent and social interaction. It is common to go to the doctor to get assistance on how to better manage our health. On the contrary, in our study, the AI had a task-oriented interaction with subjects in a nonsocial setting. As \@BBOPcitet\@BAP\@BBNYang et al., (2022)\@BBCP argue, when the interaction occurs in nonsocial scenarios, AI products are treated as utilities, resulting in an indifference toward its degree of anthropomorphism. Consequently, contrary to the belief that human-like AI identities facilitate human-AI interaction, in certain cases, designers of AI tools should not focus their efforts on anthropomorphism \@BBOPcitep\@BAP\@BBN(Yang et al.,, 2022)\@BBCP. The findings of our experiment align with this explanation, suggesting that a more neutral identity would be preferable in certain cases.
4.3 Lack of manipulation checks
An important limitation of this study, which should be addressed, is the absence of manipulation checks. In the case of AI identities, the study did not measure subjects’ perceived humanness. Consequently, it is unclear whether participants actually perceived the human-like identity as more humanlike compared to the generic or the robot-like conditions. In other words, we cannot determine if subjects anthropomorphized the AI. Therefore, the observed differences cannot be conclusively attributed to perceptions of humanness, but rather to the specific manipulations employed. For instance, differences between the generic condition and the humanlike condition may have been due to the characteristics of the image employed, rather than the degree of humanness. Similarly, the study did not measure the perceived positive or negative sentiment of the framing conditions. As mentioned above, the absence of a significant framing effect might be explained by the failure of the experimental manipulation to induce framing bias in subjects. This lack of measurement is a significant limitation of the study and should be considered.